the length of a room is four feet more than twice the width. If the perimeter is 92 feet. find the length and the width of the room.
Answer by Ruli(21)

Perimeter = 2L+2w
The length of a room is four feet more than twice the width = (2x+4)
2(2x+4) + 2(x) = 92
4x + 8 + 2x = 92
6x + 8 = 92
6x + 8 - 8 = 92 - 8
6x = 84
Divide each side by 6
x = 14
The length of the room is 32ft and the width is 14ft
check:
2[2(14)+4] + 2(14) = 92
2[28+4] + 28 = 92
2(32) + 28 = 92
64 + 28 = 92
92 = 92